OpenStudy (anonymous):

Find the sum: 20 + (-22) = A. -42 B. -2 C. 2 D. 42

OpenStudy (anonymous):

@Itrytohelpu

OpenStudy (anonymous):

20+(-22)=20-22=-2 Ansver is B

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

help other one:)